OpenStudy (anonymous):

There are a total of 78 baseball and football cards in a box. The ratio of baseball cards to football cards 8 to 5. How many baseball cards are in the box?

OpenStudy (anonymous):

8:5 so 8+5 = 13 78/13 = 6 6*8 = 48 6*5 = 30 CHECK: 30+ 48 = 78 (Its all good) Baseball cards = 48 Cards :)